OpenStudy (anonymous):

Help Please. i think it would be 9x^2√x I'm not sure though. Simplify √81x^3.

OpenStudy (anonymous):

Is x^3 inside the square root as well? Or only 81?

OpenStudy (anonymous):

It's in the square root as well.

OpenStudy (mertsj):

\[\sqrt{81x^3}=\sqrt{81}\sqrt{x^2}\sqrt{x}=9x \sqrt{x}\]

OpenStudy (anonymous):

Thank you!

OpenStudy (mertsj):

yw